Coldstream Corp. is comparing two different capital structures. Plan I would result in 12,000 shares of stock and $100,000 in debt. Plan II would result in 4,000 shares of stock and $200,000 in debt. The interest rate on the debt is 8 percent.

a. Ignoring taxes, compare both of these plans to an all-equity plan assuming that EBIT will be $70,000. The all-equity plan would result in 20,000 shares of stock outstanding. What is the EPS for each of these plans? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All equity $


b.
In part (a), what are the break-even levels of EBIT for each plan as compared to that for an all-equity plan? (Do not round intermediate calculations.)

EBIT
Plan I and all-equity $
Plan II and all-equity $


c.
Ignoring taxes, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations.)

EBIT           $

d-1
Assuming that the corporate tax rate is 40 percent, what is the EPS of the firm? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All equity $


d-2
Assuming that the corporate tax rate is 40 percent, what are the break-even levels of EBIT for each plan as compared to that for an all-equity plan? (Do not round intermediate calculations.)

EBIT
Plan I and all-equity $
Plan II and all-equity $


d-3
Assuming that the corporate tax rate is 40 percent, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations.)

EBIT           $
Answer #1
Plan I Plan II All Equity
No of Shares (i)            12,000            4,000           20,000
Equity Capital $ (12.5/ Share) (ii)          150,000         50,000        250,000
Debt (iii)          100,000       200,000                    -  
Total Investment in the business (ii+iii)          250,000       250,000        250,000
a EBIT            70,000         70,000           70,000
Interset @ 8% on debt (iii*8%)              8,000         16,000                    -  
EBT            62,000         54,000           70,000
Tax (Ignored)                     -                     -                      -  
EAT            62,000         54,000           70,000
Earning Per Share (EAT/i)                 5.17            13.50               3.50
  

b EBIT Level where EPS would be equal

Plan I and all-equity $ 20000

Plan II and all-equity $ 20000

b EBIT            20,000         20,000           20,000
Interset @ 8% on debt (iii*8%)              8,000         16,000                    -  
EBT            12,000            4,000           20,000
Tax (0%)                     -                     -                      -  
EAT            12,000            4,000           20,000
Earning Per Share (EAT/i)                 1.00              1.00               1.00
c What level of EBIT would give equal EPS for Plan 1 and II

(X-8000)/12000 = (X-16000)/4000

(X-8000)/3 = (X-16000)

X-8000 = 3X-48000

2X=40000

X = $ 20000 ( At a EBIT level of $ 20000, plan I and II would give equal EPS.

d-1 EBIT            70,000         70,000           70,000
Interset @ 8% on debt (iii*8%)              8,000         16,000                    -  
EBT            62,000         54,000           70,000
Tax (40%)            24,800         21,600           28,000
EAT            37,200         32,400           42,000
Earning Per Share (EAT/i)                 3.10              8.10               2.10

d-2 EBIT Level where EPS would be equal with Tax rate of 40%

Plan I and all-equity $ 20000

Plan II and all-equity $ 20000

d-2 EBIT            20,000         20,000           20,000
Interset @ 8% on debt (iii*8%)              8,000         16,000                    -  
EBT            12,000            4,000           20,000
Tax (40%)              4,800            1,600             8,000
EAT              7,200            2,400           12,000
Earning Per Share (EAT/i)                 0.60              0.60               0.60

d-3 EBIT Level where EPS would be equal with tax rate of 40% for Plans I and II is $ 20000

> Thank you!

Tariq Armstead Wed, Dec 1, 2021 2:30 AM